. Find the square of the following numbers.   v. 71  
j,khfmkfygkmfygjm

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


v. (71)2
= (70+1 )2
= (70)2+ (1)2+2×70×1 =4900+1+140
= 5041




Thanks